OpenStudy (anonymous):

se the distributive property to rewrite the expression and solve. 8 63 = 8 (Blank1) + 8 (Blank2) 3 = I think blank1 is 3

OpenStudy (anonymous):

8 * 63

jimthompson5910 (jim_thompson5910):

hint: 8*63 = 8*(63) = 8*(60+3)

OpenStudy (anonymous):

8 * 63 = 8 * (Blank + 8 * 3 = Blank

jimthompson5910 (jim_thompson5910):

another hint: a*(b+c) = a*b + a*c

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):